OpenStudy (anonymous):

use complete sentences to explain each step in simplifying the ratio x^2+8x/x^2+10x+16 and provide the simplified form

OpenStudy (anonymous):

X^2+8X = x(x+8) x^2+10x+16 = (x+2)(x+8) now try take ratio of both you will get your answer

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

ok. welcome..